# SoftHGNN: Soft Hypergraph Neural Networks for General Visual Recognition

### v1

@@ Abstract
I Introduction
In recent years, significant advancements in visual recognition have primarily been driven by deep neural network-based models, especially Convolutional Neural Networks (CNNs) and Transformer models [1, 2, 3, 4]. CNNs excel at capturing local spatial patterns, whereas Transformer models, built upon self-attention mechanisms, effectively model global relationships among image tokens. Current mainstream paradigms, such as Vision Transformer (ViT) [5] and its variants [6, 7, 8, 9], typically process images into sequences of visual tokens, emphasizing the modeling of pair-wise relations between these tokens.
@@ II Related Work
Hypergraph Neural Networks (HGNNs) incorporate neural networks into hypergraph representation learning. HGNNs leverage hypergraph convolution, a two‑step message‑passing scheme that first aggregates vertex features into hyperedges and then redistributes them back to vertices, to capture high‑order correlations [11, 14]. Compared to traditional Graph Neural Networks (GNNs), this high-order information aggregation mechanism significantly enhances the representational power of HGNNs[11, 20]. Feng et al. [13] proposed a spectral-domain hypergraph convolution method that demonstrated outstanding performance on high-order relational data. Building on this, Gao et al. [21] introduced a more flexible approach for modeling higher‑order relationships along with a generalized spatial hypergraph convolution framework. It extends both the adaptability of hypergraph modeling and the universality of hypergraph convolution operations. Hypergraph learning has already shown significant advantages in modeling high-order relationships across various domains such as social networks [22, 23], knowledge graphs [24], recommendation systems [25, 26], and bioinformatics [27, 28]. It is important to note, however, that these tasks are inherently based on non-Euclidean data with graph or hypergraph structures. Hypergraphs have not yet been widely applied to data in explicit Euclidean spaces, such as visual data like images.
In recent years, research has gradually begun to explore the incorporation of hypergraphs and hypergraph neural networks into visual recognition tasks. Feng et al. proposed Hyper-YOLO [15], which integrates a hypergraph computation module into the Neck component of the YOLO model [29] to achieve cross-level and cross-spatial hypergraph semantic aggregation. Han et al. introduced Vision HGNN [30], exploring the replacement of Transformer modules with hypergraph convolution modules. Wang et al. proposed HGFormer [31], which establishes many-to-many topological connections at the patch level of images and utilizes a hypergraph attention mechanism to learn global visual relationships. Chen et al. proposed HI2R [32], which simultaneously extracts the structural information within a category and the relationship information between categories through hypergraphs for more accurate fine-grained visual recognition.
@@
III Preliminaries on Traditional Hypergraph
III-A Hypergraph
@@
III-B Hypergraph Neural Networks
@@ III-C Complexity Analysis
III-C1 Complexity of Hypergraph Construction
@@
III-C2 Complexity of Hypergraph Neural Networks
@@
IV Proposed Method
IV-A Definition of Soft Hypergraph
@@
where $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ is the set of vertices, comprising $N$ vertices. =$\mathcal{E}=\{e_{1},e_{2},\ldots,e_{M}\}$ is the set of hyperedges, comprising $M$ hyperedges. Note that in fact, each hyperedge is the set of all vertices. $A\in[0,1]^{N\times M}$ is the participation matrix between vertices and hyperedges, where $A_{n,m}$ represents the degree to which vertex $v_{n}$ participates in hyperedge $e_{m}$, with its values ranging between 0 and 1.
@@
In summary, our proposed soft design offers two significant advantages. On the one hand, continuous and differentiable hyperedges can more delicately describe the relationships between vertices and hyperedges, capturing varying degrees of association. For example, in scenarios with partial occlusion or ambiguity, the relationship between vertices and hyperedges is often not absolute, and soft hypergraphs are better able to reflect such gradual and uncertain interactions. On the other hand, soft hypergraphs enable the dynamic and learnable generation of hyperedges, allowing the model to adaptively capture complex patterns hidden in large-scale data. This characteristic is crucial for constructing efficient and flexible models in practical visual tasks (a detailed discussion will be presented in Section IV-C).
IV-B Overview of SoftHGNN
@@
IV-C Soft Hyperedge Generation
@@
IV-C1 Dynamic Hyperedge Prototype Generation
@@
IV-C2 Vertex Feature Pre-projection
@@
$X_{\text{proj}}=W_{\text{pre}}X\in\mathbb{R}^{N\times D},$ (6)
@@
IV-C3 Participation Matrix Generation
@@
Here, dividing by $\sqrt{D_{\text{head}}}$ stabilizes the gradients by preventing the dot-product values from growing too large as the dimensionality increases. To fuse the information across all heads, we average the similarity scores: $S=\frac{1}{h}\sum_{\tau=1}^{h}S^{(\tau)}\in\mathbb{R}^{N\times M}$. Finally, we apply the Softmax function to normalize the similarity vector for each vertex, yielding the continuous and differentiable participation matrix:
$A_{i,j}=\frac{\exp(S_{i,j})}{\sum_{k=1}^{M}\exp(S_{i,k})},$ (8)
@@
IV-D Message Passing on Soft Hypergraph
The SoftHGNN leverage the soft hyperedge structure to enable effective interactions between vertices. The general message passing is divided into two stages: Aggregation from Vertices to Soft Hyperedges and Dissemination from Soft Hyperedges to Vertices.
IV-D1 Aggregation from Vertices to Soft Hyperedges
@@
IV-D2 Dissemination from Soft Hyperedges to Vertices
@@
IV-D3 Matrix Formulation of Message Passing
@@
IV-E Sparse Soft Hyperedge Selection and Load-Balancing
@@
IV-E1 Sparse Hyperedge Selection Strategy
@@
$\begin{gathered}\mathcal{E}=\mathcal{E}^{\mathrm{fixed}}\cup\mathcal{E}^{\mathrm{dyn}},\quad\bigl{|}\mathcal{E}^{\mathrm{fixed}}\bigr{|}=M_{\mathrm{fixed}},\quad\bigl{|}\mathcal{E}^{\mathrm{dyn}}\bigr{|}=M_{\mathrm{dyn}},\\ \mathrm{s.t.\ }M=M_{\mathrm{fixed}}+M_{\mathrm{dyn}}.\end{gathered}$ (16)
@@
where $S^{\text{dyn}}$ denotes the portion of the initial participation matrix $S$ corresponding to the $M_{\text{dyn}}$ dynamic soft hyperedges. For the dynamic soft hyperedges, we select only the top-$k$ hyperedges with the highest activation scores. Theoretically, this step can be viewed as selecting a subset $\mathcal{E}^{\text{sel}}\subset\mathcal{E}^{\text{dyn}}$ with $|\mathcal{E}^{\text{sel}}|=k$. Finally, the hyperedge set is updated as $\mathcal{E}=\mathcal{E}^{\text{fixed}}\cup\mathcal{E}^{\text{sel}},$ and the corresponding updated unnormalized participation matrix becomes $S^{\prime}\in\mathbb{R}^{N\times(M_{\text{fixed}}+k)}.$ After applying the Softmax function to each vertex’s assignment vector, we obtain the final continuous and differentiable participation matrix used for message passing:
$A_{i,j}^{\text{total}}=\frac{\exp(S_{i,j}^{\prime})}{\sum_{q=1}^{M_{\text{fixed}}+k}\exp(S_{i,q}^{\prime})},$ (18)
@@
IV-E2 Load-Balancing Regularizer
@@
$\mathcal{M}_{j}^{(t)}=\begin{cases}1,&\text{if dynamic soft hyperedge }e_{j}\text{ is selected},\\ 0,&\text{otherwise}.\end{cases}$ (19)
@@
This regularization term encourages the dynamic soft hyperedges to be selected uniformly over multiple forward passes, ensuring that each dynamic hyperedge contributes meaningfully to the message passing process rather than relying solely on a few hyperedges, thereby enhancing the model’s robustness and the effective utilization of hyperedges.
IV-F Complexity Analysis of SoftHGNN
@@
IV-F1 The Stage of Soft Hyperedge Generation
@@
IV-F2 The Stage of Message Passing
@@ IV-F3 Overall Discussion
V Experimental Results
V-A Tasks and Datasets
@@ V-B Evaluation Metrics
TABLE I: Classification accuracy on CIFAR-10 and CIFAR-100 datasets.
Method CIFAR-10 ACC (%) CIFAR-100 ACC (%)
@@
+ SoftHGNN (Ours) 80.5 ($\uparrow$1.9) 55.2 ($\uparrow$1.8)
+ SoftHGNN-SeS (Ours) 81.1 ($\uparrow$2.5) 55.8 ($\uparrow$2.4)
V-C Experimental Setup
@@
By default, both the number of heads and the number of soft hyperedges in SoftHGNN are set to 8. In the variant with Soft Hyperedge Selection, referred to as SoftHGNN-SeS, the number of fixed soft hyperedges is set to 16, with an additional 32 dynamic soft hyperedges from which a sparse subset is selected. A sparsity ratio of 50% is applied, meaning that 16 out of the 32 dynamic soft hyperedges are activated during each forward pass. When using SoftHGNN-SeS, the final loss is computed as the sum of the load-balancing regularization loss and the original loss of the baseline model.
[b] Method Venue Backbone SHTech-Part A SHTech-Part B Param. (M) MAE ($\downarrow$) MSE ($\downarrow$) MAE ($\downarrow$) MSE ($\downarrow$) CSRNet [41] CVPR-2018 VGG-16 [42] 68.2 115.0 10.6 16.0 16.3 BL [43] ICCV-2019 VGG-19 [42] 62.8 101.8 7.7 12.7 21.5 DM-Count [44] NeurIPS-2020 VGG-19 [42] 59.7 95.7 7.4 11.8 21.5 TopoCount [45] AAAI-2021 VGG-16 [45] 61.2 104.6 7.8 13.7 25.8 LSC-CNN [37] TPAMI-2021 VGG-16 [42] 66.4 117.0 8.1 12.7 35.1 Ctrans-MISN [46] PRAI-2022 Swin-Trans [8] 55.8 95.9 7.3 11.4 - CLTR [47] ECCV-2022 ViT [5] 56.9 95.2 6.5 10.6 43.5 GMS [48] TIP-2023 HRNet [49] 68.8 138.6 16.0 33.5 - GCFL [50] TPAMI-2023 VGG-19 [42] 57.5 94.3 6.9 11.0 21.5 Gramformer [51] AAAI-2024 VGG-19 [42] 54.7 87.1 - - 29.1 CAAPN [36] TPAMI-2024 VGG-16 [42] 54.4 97.3 5.8 9.8 - mPrompt [52] CVPR-2024 HRNet [49] 53.2 85.4 6.3 9.8 48.3 CCTrans [35] arXiv-2021 Twins-PCPVT [9] 54.8 89.0 7.6 12.2 106.4 + Self-Attention [5] - Twins-PCPVT [9] 54.7 ($\downarrow$0.2%) 89.7 ($\uparrow$0.8%) 7.4 ($\downarrow$2.6%) 10.9 ($\downarrow$10.7%) 107.6 + HGNN [21] - Twins-PCPVT [9] 53.7 ($\downarrow$2.0%) 82.7 ($\downarrow$7.1%) 7.2 ($\downarrow$5.3%) 10.9 ($\downarrow$10.7%) 108.0 + SoftHGNN (Ours) - Twins-PCPVT [9] 52.8 ($\downarrow$3.6%) 82.1 ($\downarrow$7.8%) 7.1 ($\downarrow$6.6%) 10.6 ($\downarrow$13.1%) 107.6 + SoftHGNN-SeS (Ours) - Twins-PCPVT [9] 53.4 ($\downarrow$2.6%) 82.0 ($\downarrow$7.9%) 7.0 ($\downarrow$7.9%) 10.6 ($\downarrow$13.1%) 112.9 CCTrans [35] arXiv-2021 Pyramid ViT [7] 53.7 85.9 7.4 12.5 49.9 + Self-Attention [5] - Pyramid ViT [7] 53.1 ($\downarrow$1.1%) 83.1 ($\downarrow$3.3%) 7.1 ($\downarrow$4.1%) 10.9 ($\downarrow$12.8%) 51.2 + HGNN [21] - Pyramid ViT [7] 53.5 ($\downarrow$0.4%) 85.5 ($\downarrow$0.5%) 7.0 ($\downarrow$5.4%) 10.9 ($\downarrow$12.8%) 51.6 + SoftHGNN (Ours) - Pyramid ViT [7] 51.7 ($\downarrow$3.7%) 81.3 ($\downarrow$5.4%) 6.8 ($\downarrow$8.1%) 10.7 ($\downarrow$14.4%) 51.2 + SoftHGNN-SeS (Ours) - Pyramid ViT [7] 51.7 ($\downarrow$3.7%) 79.2 ($\downarrow$7.8%) 6.6 ($\downarrow$10.8%) 10.5 ($\downarrow$16.0%) 56.4
TABLE II: Quantitative result of object counting performance on ShanghaiTech Part-A and Part-B datasets.
V-D Quantitative Results
V-D1 Image Classification

### v2

@@ Abstract
1 Introduction
In recent years, significant advancements in visual recognition have primarily been driven by deep neural network-based models, especially Convolutional Neural Networks (CNNs) and Transformer models [1, 2, 3]. CNNs excel at capturing local spatial patterns, whereas Transformer models, built upon self-attention mechanisms, effectively model global relationships among image tokens. Current mainstream paradigms, such as Vision Transformer (ViT) [4] and its variants [5, 6, 7, 8, 9], typically process images into sequences of visual tokens, emphasizing the modeling of pair-wise relations between these tokens.
@@ 2 Related Work
Hypergraph Neural Networks (HGNNs) incorporate neural networks into hypergraph representation learning. HGNNs leverage hypergraph convolution, a two‑step message‑passing scheme that first aggregates vertex features into hyperedges and then redistributes them back to vertices, to capture high‑order correlations [11, 14, 20]. Compared to traditional Graph Neural Networks (GNNs), this high-order information aggregation mechanism significantly enhances the representational power of HGNNs[11, 21]. Feng et al. [13] proposed a spectral-domain hypergraph convolution method that demonstrated outstanding performance on high-order relational data. Building on this, Gao et al. [22] introduced a more flexible approach for modeling higher‑order relationships along with a generalized spatial hypergraph convolution framework. It extends both the adaptability of hypergraph modeling and the universality of hypergraph convolution operations. Hypergraph learning has already shown significant advantages in modeling high-order relationships across various domains such as social networks [23, 24], knowledge graphs [25], recommendation systems [26, 27], action recognition [28, 29], and bioinformatics [30, 31]. It is important to note, however, that these tasks are inherently based on non-Euclidean data with graph or hypergraph structures. Hypergraphs have not yet been widely applied to data in explicit Euclidean spaces, such as visual data like images.
In recent years, research has gradually begun to explore the incorporation of hypergraphs and hypergraph neural networks into visual recognition tasks [28, 29, 32]. Feng et al. proposed Hyper-YOLO [15], which integrates a hypergraph computation module into the Neck component of the YOLO model [33] to achieve cross-level and cross-spatial hypergraph semantic aggregation. Han et al. introduced Vision HGNN [34], exploring the replacement of Transformer modules with hypergraph convolution modules. Wang et al. proposed HGFormer [35], which establishes many-to-many topological connections at the patch level of images and utilizes a hypergraph attention mechanism to learn global visual relationships. Chen et al. proposed HI2R [36], which simultaneously extracts the structural information within a category and the relationship information between categories through hypergraphs for more accurate fine-grained visual recognition.
@@
3 Preliminaries on Traditional Hypergraph
3.1 Hypergraph
@@
3.2 Hypergraph Neural Networks
@@ 3.3 Complexity Analysis
3.3.1 Complexity of Hypergraph Construction
@@
3.3.2 Complexity of Hypergraph Neural Networks
@@
4 Proposed Method
4.1 Definition of Soft Hypergraph
@@
where $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ is the set of vertices, comprising $N$ vertices. $\mathcal{E}=\{e_{1},e_{2},\ldots,e_{M}\}$ is the set of hyperedges, comprising $M$ hyperedges. Note that in fact, each hyperedge is the set of all vertices. $A\in[0,1]^{N\times M}$ is the participation matrix between vertices and hyperedges, where $A_{n,m}$ represents the degree to which vertex $v_{n}$ participates in hyperedge $e_{m}$, with its values ranging between 0 and 1.
@@
In summary, our proposed soft design offers two significant advantages. On the one hand, continuous and differentiable hyperedges can more delicately describe the relationships between vertices and hyperedges, capturing varying degrees of association. For example, in scenarios with partial occlusion or ambiguity, the relationship between vertices and hyperedges is often not absolute, and soft hypergraphs are better able to reflect such gradual and uncertain interactions. On the other hand, soft hypergraphs enable the dynamic and learnable generation of hyperedges, allowing the model to adaptively capture complex patterns hidden in large-scale data. This characteristic is crucial for constructing efficient and flexible models in practical visual tasks (a detailed discussion will be presented in Section 4.3).
4.2 Overview of SoftHGNN
@@
4.3 Soft Hyperedge Generation
@@
4.3.1 Dynamic Hyperedge Prototype Generation
@@
4.3.2 Vertex Feature Pre-projection
@@
$X_{\text{proj}}=XW_{\text{pre}}\in\mathbb{R}^{N\times D},$ (6)
@@
4.3.3 Participation Matrix Generation
@@
Here, dividing by $\sqrt{D_{\text{head}}}$ stabilizes the gradients by preventing the dot-product values from growing too large as the dimensionality increases. To fuse the information across all heads, we average the similarity scores: $S=\frac{1}{h}\sum_{\tau=1}^{h}S^{(\tau)}\in\mathbb{R}^{N\times M}$. Finally, we apply Softmax along the node dimension for each hyperedge’s node assignment vector, yielding the continuous and differentiable participation matrix:
$A_{i,j}=\frac{\exp(S_{i,j})}{\sum_{k=1}^{N}\exp(S_{k,j})},$ (8)
@@
4.4 Message Passing on Soft Hypergraph
The SoftHGNN leverages the soft hyperedge structure to enable effective interactions between vertices. The general message passing is divided into two stages: Aggregation from Vertices to Soft Hyperedges and Dissemination from Soft Hyperedges to Vertices.
4.4.1 Aggregation from Vertices to Soft Hyperedges
@@
4.4.2 Dissemination from Soft Hyperedges to Vertices
@@
4.4.3 Matrix Formulation of Message Passing
@@
4.5 Sparse Soft Hyperedge Selection and Load-Balancing
@@
4.5.1 Sparse Hyperedge Selection Strategy
@@
$\small\begin{gathered}\mathcal{E}=\mathcal{E}^{\mathrm{fixed}}\cup\mathcal{E}^{\mathrm{dyn}},~\bigl{|}\mathcal{E}^{\mathrm{fixed}}\bigr{|}=M_{\mathrm{fixed}},~\bigl{|}\mathcal{E}^{\mathrm{dyn}}\bigr{|}=M_{\mathrm{dyn}},\\ \mathrm{s.t.\ }M=M_{\mathrm{fixed}}+M_{\mathrm{dyn}}.\end{gathered}$ (16)
@@
where $S^{\text{dyn}}$ denotes the portion of the initial participation matrix $S$ corresponding to the $M_{\text{dyn}}$ dynamic soft hyperedges. For the dynamic soft hyperedges, we select only the top-$k$ hyperedges with the highest activation scores. Theoretically, this step can be viewed as selecting a subset $\mathcal{E}^{\text{sel}}\subset\mathcal{E}^{\text{dyn}}$ with $|\mathcal{E}^{\text{sel}}|=k$. Finally, the hyperedge set is updated as $\mathcal{E}=\mathcal{E}^{\text{fixed}}\cup\mathcal{E}^{\text{sel}},$ and the corresponding updated unnormalized participation matrix becomes $S^{\prime}\in\mathbb{R}^{N\times(M_{\text{fixed}}+k)}.$ After applying Softmax to each hyperedge’s node assignment vector, we obtain the final continuous and differentiable participation matrix used for message passing:
$A_{i,j}^{\text{total}}=\frac{\exp(S_{i,j}^{\prime})}{\sum_{q=1}^{N}\exp(S_{q,j}^{\prime})},$ (18)
@@
4.5.2 Load-Balancing Regularizer
@@
$\mathcal{M}_{j}^{(t)}=\begin{cases}1,\text{if dynamic soft hyperedge }e_{j}\text{ selected},\\ 0,\text{otherwise}.\end{cases}$ (19)
@@
This regularization term encourages the dynamic soft hyperedges to be selected uniformly over multiple forward passes, ensuring that each dynamic hyperedge contributes meaningfully to the message passing process rather than relying solely on a few hyperedges, thereby enhancing the model’s robustness as well as the effective utilization of hyperedges.
4.6 Complexity Analysis of SoftHGNN
@@
4.6.1 The Stage of Soft Hyperedge Generation
@@
4.6.2 The Stage of Message Passing
@@ 4.6.3 Overall Discussion
5 Experimental Results
5.1 Tasks and Datasets
@@ 5.2 Evaluation Metrics
Table 1: Classification accuracy (%) on CIFAR-10 and CIFAR-100 datasets.
Method CIFAR-10 CIFAR-100
@@
+ SoftHGNN 80.5 ($\uparrow$1.9) 55.2 ($\uparrow$1.8)
+ SoftHGNN-SeS 81.1 ($\uparrow$2.5) 55.8 ($\uparrow$2.4)
5.3 Experimental Setup
Table 2: Quantitative result of counting performance on ShanghaiTech Part-A and Part-B datasets.
[b]
Method Backbone SHTech-Part A SHTech-Part B Param.
MAE ($\downarrow$) MSE ($\downarrow$) MAE ($\downarrow$) MSE ($\downarrow$)
CSRNet [42] VGG-16 [43] 68.2 115.0 10.6 16.0 16.3 M
BL [44] VGG-19 [43] 62.8 101.8 7.7 12.7 21.5 M
DM-Count [45] VGG-19 [43] 59.7 95.7 7.4 11.8 21.5 M
TopoCount [46] VGG-16 [46] 61.2 104.6 7.8 13.7 25.8 M
LSC-CNN [41] VGG-16 [43] 66.4 117.0 8.1 12.7 35.1 M
Ctrans-MISN [47] Swin-Trans [6] 55.8 95.9 7.3 11.4 -
CLTR [48] ViT [4] 56.9 95.2 6.5 10.6 43.5 M
GMS [49] HRNet [50] 68.8 138.6 16.0 33.5 -
GCFL [51] VGG-19 [43] 57.5 94.3 6.9 11.0 21.5 M
Gramformer [52] VGG-19 [43] 54.7 87.1 - - 29.1 M
CAAPN [40] VGG-16 [43] 54.4 97.3 5.8 9.8 -
mPrompt [53] HRNet [50] 53.2 85.4 6.3 9.8 48.3 M
CCTrans [39] Twins-PCPVT [54] 54.8 89.0 7.6 12.2 106.4 M
+ Self-Attn [4] Twins-PCPVT [54] 54.7 ($\downarrow$0.2%) 89.7 ($\uparrow$0.8%) 7.4 ($\downarrow$2.6%) 10.9 ($\downarrow$10.7%) 107.6 M
+ HGNN [22] Twins-PCPVT [54] 53.7 ($\downarrow$2.0%) 82.7 ($\downarrow$7.1%) 7.2 ($\downarrow$5.3%) 10.9 ($\downarrow$10.7%) 108.0 M
+ SoftHGNN Twins-PCPVT [54] 52.8 ($\downarrow$3.6%) 82.1 ($\downarrow$7.8%) 7.1 ($\downarrow$6.6%) 10.6 ($\downarrow$13.1%) 107.6 M
+ SoftHGNN-SeS Twins-PCPVT [54] 53.4 ($\downarrow$2.6%) 82.0 ($\downarrow$7.9%) 7.0 ($\downarrow$7.9%) 10.6 ($\downarrow$13.1%) 112.9 M
CCTrans [39] Pyramid ViT [55] 53.7 85.9 7.4 12.5 49.9 M
+ Self-Attn [4] Pyramid ViT [55] 53.1 ($\downarrow$1.1%) 83.1 ($\downarrow$3.3%) 7.1 ($\downarrow$4.1%) 10.9 ($\downarrow$12.8%) 51.2 M
+ HGNN [22] Pyramid ViT [55] 53.5 ($\downarrow$0.4%) 85.5 ($\downarrow$0.5%) 7.0 ($\downarrow$5.4%) 10.9 ($\downarrow$12.8%) 51.6 M
+ SoftHGNN Pyramid ViT [55] 51.7 ($\downarrow$3.7%) 81.3 ($\downarrow$5.4%) 6.8 ($\downarrow$8.1%) 10.7 ($\downarrow$14.4%) 51.2 M
+ SoftHGNN-SeS Pyramid ViT [55] 51.7 ($\downarrow$3.7%) 79.2 ($\downarrow$7.8%) 6.6 ($\downarrow$10.8%) 10.5 ($\downarrow$16.0%) 56.4 M
@@ 5.3 Experimental Setup
By default, both the number of heads and the number of soft hyperedges in SoftHGNN are set to 8. In the variant with Soft Hyperedge Selection, referred to as SoftHGNN-SeS, the number of fixed soft hyperedges is set to 16, with an additional 32 dynamic soft hyperedges from which a sparse subset is selected. A sparsity ratio of 50% is applied, meaning that 16 out of the 32 dynamic soft hyperedges are activated during each forward pass. When using SoftHGNN-SeS, the final loss is computed as the sum of load-balancing regularization loss and the original loss of the baseline.
Table 3: Quantitative results of nano size model object detection performance on MS COCO dataset.
[b] Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param. Gold-YOLO-N [59] 55.7 39.6 12.1 G 2.92 ms 5.6 M Hyper-YOLO-T [15] 54.5 38.5 9.6 G 2.5 ms 3.1 M YOLOv8-N [60] 52.6 37.4 8.7 G 1.77 ms 3.2 M YOLOv9-T [61] 53.1 38.3 7.7 G 2.48 ms 2.0 M YOLOv10-N [62] 53.8 38.5 6.7 G 1.84 ms 2.3 M YOLO11-N [57] 55.3 39.4 6.5 G 1.50 ms 2.6 M YOLO11-N [57] 53.6 38.2 6.5 G 1.59 ms 2.6 M YOLO11-SoftHGNN-N 55.8 ($\uparrow$2.2) 40.2 ($\uparrow$2.0) 8.0 G 1.99 ms 3.2 M YOLOv12-N [58] 55.4 40.1 6.0 G 1.60 ms 2.5 M YOLOv12-N [58] 55.1 39.6 6.0 G 1.63 ms 2.5 M YOLOv12-SoftHGNN-N 57.0 ($\uparrow$1.9) 41.3 ($\uparrow$1.7) 7.4 G 1.99 ms 3.1 M
5.4 Quantitative Results
5.4.1 Image Classification